OpenStudy (anonymous):

Need a little help proving my answer - Generate a function for the following data: x = -10, -8, -6, -4, -2, 0, 2, 4, 6, 8, 10. y = -5, -4, -3, -2, -1, 0, -1, -2, -3, -4, -5. - See comment for a better view, my answer is included.

OpenStudy (anonymous):

\[x = -10, -8, -6, -4, -2, 0, 2, 4, 6, 8, 10\]\[y = -5, -4, -3, -2, -1, 0, 1, 2, 3, 4, 5\] My answer: \[f(x) = \frac{ -(\sqrt{x ^{2}}) }{ 2 }\] Just wondering if anybody got the same answer as I.

OpenStudy (anonymous):

Wonder why \(-\sqrt x^2\)?

OpenStudy (anonymous):

but x/2?

OpenStudy (michele_laino):

please note that: \[\sqrt{x ^{2}}=\pm x\]

OpenStudy (michele_laino):

I think, as I read from your data, that we can write: \[x=2y\]

OpenStudy (anonymous):

I need to make x a negative all the time so that y is always a negative.

OpenStudy (anonymous):

oh, typo again

OpenStudy (anonymous):

-10/2=-5 right?

OpenStudy (anonymous):

y = are all negative

OpenStudy (anonymous):

The blue question box has the right data.

OpenStudy (anonymous):

How about \(y=\dfrac{-|x|}{2}\)?

OpenStudy (anonymous):

Oh, nice idea there! Just waiting for Mich as a double check.

OpenStudy (michele_laino):

with your formula I'm not able to get positive y's @OOOPS

OpenStudy (anonymous):

@Michele_Laino he edited his problem.!!

OpenStudy (anonymous):

@OOOPS My bad

OpenStudy (anonymous):

but doesn't |x| absolute? (-1)(|x|) doesn't make a negative?

OpenStudy (michele_laino):

Sorry, but from the text I see y=0, 1, 2, 3,... @OOOPS

OpenStudy (anonymous):

@Michele_Laino He said: all y's are negative!!

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

sorry again! @OOOPS

OpenStudy (anonymous):

@Sadworld |x| make the value of x positive and (-1) make value of y is negative.

OpenStudy (anonymous):

So you're function is a possible answer to the problem? Mich, may you double check? I get the same answer as OOOPS